OpenStudy (anonymous):

Complete the ordered pair so that it is a solution of x + y - 1 = 0. (-6, ___) 3 4 5

OpenStudy (anonymous):

1. (-6, 5) 2. (4/3, 0) 3. y = x/2 + 6 4. y = 3/2x -5/2 5. slope = -6/5 6. slope = 0 7. shift 6 units right